OpenStudy (anonymous):

If co2 and oxygen are mixed in the mole ratio 1:9 , the density (g/L) of the gas at stp is?

mathslover (mathslover):

mass : volume = density

mathslover (mathslover):

\[\huge{density=\frac{mass}{volume}}\]

OpenStudy (anonymous):

So how to connet these

mathslover (mathslover):

22.4 L ---> 1 mole at STP 1 mole means we can take its molecular mass

mathslover (mathslover):

calculate the volume of CO2 and O2 :::: (22.4 L * 1 ) + (22.4 L * 9) = 224 L Calculate the mass of CO2 and O2 : CO2 = 12 + 16(2) = 44 g. O2 = 16*2 = 32 g. can u calculate the density now ?

OpenStudy (anonymous):

sorry it is oxygen not o2

OpenStudy (anonymous):

Plz read the question Now...

OpenStudy (anonymous):

and sorry the asnwer should be 1.5

mathslover (mathslover):

i think that should be o2 ... as i am getting 1.5 if that is o2 ..

OpenStudy (anonymous):

ok then plz show hw u got it...

mathslover (mathslover):

for oxygen and not o2 we have : mass = 44g + 9(16)g = 44g + 144g= 188g and total volume : 224 L ( it will remain same ) 188/224 is not equal to 1.5

mathslover (mathslover):

But for O2

OpenStudy (anonymous):

hey @mathslover is it 22.4L

OpenStudy (anonymous):

But u have written 224L.............why?

mathslover (mathslover):

we have total mass : 44g + 9(32)g =44g + 288g = 332 g total volume : 10(22.4L) = 224 L density = 332/224 = 1.5..

OpenStudy (anonymous):

yup got it thxxx

mathslover (mathslover):

@Yahoo! we have supposed to have total no. of moles = 10 hence 1 mole of co2 and 9 moles of o2 = 1(22.4)+9(22.4)=22.4(1+9)=22.4*10 = 224 L

mathslover (mathslover):

nice to see that u have got that.... gr8